# QUIJOTE-TFGI polarization calibration Ground characterization and on-sky validation with Tau A and the Moon

### v1

@@ Abstract
Context. Polarization measurements in the 30–40 GHz range are essential to improve the understanding of Galactic synchrotron radiation and potential anomalous microwave emissions. Ultimately, this advancement will aid in detecting primordial B modes. To accomplish this, it is crucial to achieve unprecedented performance in instrumentation. Specifically, accurately calibrating instrument responses is an essential step in state-of-the-art setups. The modeling and mitigation of systematic effects are critical in this context.
@@
Results. By creating calibration datasets obtained with the reference diode, we evaluate the data quality and quantify phase switch errors to account for the fine polarization response. We also utilize Tau A and Moon observations to calibrate the system’s response and stability over time. In addition, we calculate the refraction index of the Moon to be $n_{\rm{Moon}}=1.209\pm 0.007\,\text{(stat)}\,\pm 0.005\,\text{(sys)}$ at 31 GHz under smooth-surface assumption.
@@ 1 Introduction
With the completion of these groundbreaking observations, attention has shifted towards investigating the polarization of the CMB (Abazajian et al. 2016). Despite its signal being orders of magnitude fainter than total intensity, advancements in instrumentation now permit the study of polar signals with unprecedented sensitivity. The polarization pattern of the CMB, typically decomposed into E-modes and B-modes (Seljak & Zaldarriaga 1997; Kamionkowski et al. 1997), carries critical information about the early Universe’s conditions, including the presence of primordial gravitational waves generated during inflation.While satellite-based experiments are required to measure the largest angular scales, ground-based observations have also made significant contributions to the field, especially at small angular scales where satellite telescopes struggle to achieve the necessary angular resolution. The international thrust has deployed in the last decade, among others, the Atacama B-mode Search experiment (ABS; Simon et al. 2016; Kusaka et al. 2018), the Atacama Cosmology Telescope (ACT; Li et al. 2021; Qu et al. 2024), the BICEP/Keck Array program (Hui et al. 2018; BICEP/Keck Collaboration et al. 2022), the Cosmology Large Angular Scale Surveyor (CLASS; Essinger-Hileman et al. 2014; Eimer et al. 2024), the GroundBIRD telescope (Tsujii et al. 2024), the Polarbear experiment (Inoue et al. 2016; Adachi et al. 2022, the predecessor of the Simons Array experiment; Stebor et al. 2016), the South Pole Telescope (SPT; Chown et al. 2018; Sobrin et al. 2022), Spider (Ade et al. 2022; Shaw et al. 2024), and the Q & U Bolometric Interferometer for Cosmology (QUBIC; Mennella et al. 2024).
In this context, the Q-U-I JOint TEnerife (QUIJOTE) CMB experiment aims to achieve two primary scientific objectives: firstly, to measure the tensor-to-scalar ratio (r) if larger than $\sim$0.05 (Rubiño-Martín et al. 2017); and secondly, to characterize polarized foregrounds (Betoule et al. 2009), particularly at low frequencies, shedding light on the synchrotron and anomalous microwave emissions. The current best upper limit on $r$ is $r<0.032$, at a 95% confidence level (Tristram et al. 2022). Current-generation experiments like the Simons Observatory small aperture telescopes (SO-SATs) aim at constraining $r$ with $\sigma(r)=0.003$ or better (Ade et al. 2019), while next-generation telescopes like LiteBIRD will attempt a $r>0.003$ detection with $\sigma(r)\sim 0.001$.
@@
In addition to polarization angle calibration, an equally crucial aspect is the precise calibration of detector responsivity. Any uncertainty in the gain directly affects the amplitude of the measured Stokes parameters, thereby introducing a multiplicative bias in the polarization power spectra. For next-generation CMB experiments aiming to detect primordial B-modes at levels of $r<10^{-3}$, the required responsivity calibration must be at the sub-percent level, typically better than 0.5%, to avoid systematic errors that could bias the signal (Abitbol et al. 2021). Responsivity mismatches across detectors can also induce leakage from temperature anisotropies into polarization. Precise calibration between frequency channels is also required to ensure a reliable component separation and then minimize contamination on the measurement of the cosmological signal from foreground residuals (Carralot et al. 2025).
In this work, we describe the calibration of the TFGI system by characterizing its central detector, Pixel 23 at 31 GHz. This detector is the one located in the center of the focal plane, and we use it as a benchmark for the full focal-plane calibration. In Sect. 2, we provide an overview of the TFGI instrument. Section 3 details the ground measurement, including its specific setup and mathematical framework (Sect. 3.1), data structure (Sect. 3.2) for polarization efficiency (Sect. 3.3), responsivity (Sect. 3.4), polarization angle (Sect. 3.5), and phase error (Sect. 3.6) characterization. Section 4 presents the results from an on-sky calibration and response stability using Tau A (Sect. 4.1) and the Moon (Sect. 4.2). Section 5 provides a comparison of the results across different setups. In addition, we calculate the refraction index of the Moon in Appendix A. Finally, the conclusions are presented in Sect. 6.
@@ 2 The QUIJOTE-TFGI instrument
After preliminary commissioning phases between 2018 and 2019, the TFGI was installed at the second QUIJOTE telescope (QT2) and underwent its commissioning from November 2021 to October 2022 with 7 pixels, 4 at 31 GHz and 3 at 41 GHz. The data analyzed in this paper correspond to this latter period.
@@ 3.1 Setup
The calibration process involves positioning the so-called multi-frequency calibration instrument (MFCI) in front of each detector and acquiring calibration data, as illustrated in Fig. 2. The MFCI consists of two feedhorns and OMTs, similar to the cryogenically cooled feedhorns and OMTs of the TFGI, and is connected back-to-back through two identical -20 dB waveguide couplers. The calibration signal is introduced through one coupler, while the other is terminated with a room-temperature load. The horns are aligned through a concentric stepper motor axial drive, and this setup is designed to be attached to a detector-centered grid known as a fixed plate. Additionally, a tilted aluminum flat mirror points the horn arrangement toward the sky when the cryostat is mounted on the focal plane of the QT2. In this way, the background radiation during a calibration run comes from the sky. The injected signal corresponds to the diode’s equivalent temperature when ON with a 20 dB coupling ratio. A wideband noise diode (Keysight 346CK01111https://www.keysight.com/us/en/product/346CK01/noise-source-1-ghz-50-ghz.html) was selected for its coverage of both FGI and TGI bands.
For installation, the MFCI is preassembled horizontally using a digital inclinometer and mounted onto the telescope focal plane with 0.2 deg repeatability. A dedicated acquisition system controls the stepper motor, the diode, and the phase switches. Polarization angles are referenced with respect to the diode’s E-field and defined as positive for anticlockwise rotation from the sky toward the instrument. The alignment plate is secured over the cryostat focal plane, and the MFCI is attached with spring-loaded clamps, ensuring the mirror maintains its skyward orientation. Electrical connections for the motor, diode, and switches are routed through dedicated cabling.
@@ 3.1.1 Process of polarization characterization and mathematical framework
where V is the output in volts [^2] Not to be mistaken this time with the Stokes $V$ parameter., $g$ is the gain, $\eta$ is the polarization efficiency, $\delta$ is the phase state of each polarization state (0, 90, 180, and 270 deg), and $\epsilon$ is the correction angle that represents the deviation from the nominal phase value assumed constant over the band. The subscript indicates the channel, with a comma distinguishing between TGI and FGI. Also $g$, $\epsilon$, and $\phi_{\rm{c}}$ differ between TGI and FGI. Operationally, we utilize Eq. (2c) for all channels in the on-sky analysis. All equations are, in fact, equivalent to a successive 45-degree rotation.
@@ 3.2.1 Data assessment with diode reference
[…]
The data present a linear regression slope of a few hundred nV/sample; $216\pm 113$ nV/sample in the diode ON case and $328\pm 35$ nV/sample in the diode OFF case. The uncertainties are calculated as the standard deviation of the counts. This leads to a drift in the signal smaller than 2 mV in a minute. Equivalent to $\sim$50 mK by applying the average on-sky responsivity reported in Sect. 5.2. This instability is likely due to a combination of gain drift and the instability of the optical load, originating from the background and the diode. Although this small drift is detectable, it has a negligible impact on the results discussed here.
@@ 3.4 Responsivity calibration
The responsivity calculated from the diode is primarily affected by the uncertainty in the diode’s emitted power as seen by the detector, and the fact that this measurement does not reflect the actual observational configuration used on-sky – for instance, it does not include coupling through the QT2 mirrors, which alters the beam pattern and effective system gain. While the detector’s intrinsic responsivity should, in principle, be independent of the optical setup, the total optical coupling efficiency and beam gain introduced by the optics do influence the system-level responsivity. Additionally, the relatively high temperature of the diode may induce nonlinearities in the detector response, further complicating the calibration.
@@ 3.6 Phase errors calibration
The initial fitting step aims to constrain the $\phi_{c}$ parameter for each channel. We fit Eq. (2c) under the assumption of $\epsilon=0$ deg. Consequently, the resulting $\phi_{c}$ values encompass the average $\epsilon$ across different phase angles. Subsequently, we perform the fitting for each dataset, with 19, 15, and 15 iterations, respectively, for the three datasets. We exploit the results reported in Table 1 (last row) as a hint for the initial guess of the fitting for $\phi_{c}$. The results are reported in Table 5.
@@
Table 3: Phase error corrections ($\epsilon$). [^6] The results are presented for each channel (CH) and phase state ($\delta$). The values represent averages across the datasets used, with the uncertainties calculated as the standard deviation among the different datasets. The final row reports the mean values for each phase state, with uncertainties obtained by propagating the uncertainties in the mean calculation across channels, based on the expectation that the errors should be consistent across channels by design.
@@
The results are compatible with the null value at $2\sigma$, where the $\sigma=0.65$ deg refers to the standard deviation of the uncertainties of mean values (last line) of Table 6, with a mean value of 1.3 deg. Therefore, we do not need to consider any corrections for scientific purposes at a $\sim$1-degree precision for the polarization angle. The datasets exploited for the $\epsilon$ calculation are recorded with a separation of $\sim$18 hours. The ambient variables, particularly temperature, may have influenced the response of the phase switches. More datasets at lower timescale differences might be helpful for future development. Appendix E presents the results per dataset.
@@ 4.1.1 Polarization efficiency with Tau A
For this analysis, we selected the Tau A observation with the best noise properties while being close in time to the Moon observation that will be discussed in Sect. 4.2. This observation consists of a single raster scan, and its properties are summarized in Table 7. From this observation, we produce maps of the three Stokes parameters for each channel independently, which are shown in Fig. 6. The detection of a polarized signal from Tau A is immediately clear. To estimate the source flux density, we use a beam-fitting photometry that involves fitting the $I$, $Q$, and $U$ maps of each channel to a 2D Gaussian with fixed position (using the nominal coordinates of Tau A) and width of 1 deg. This fit then involves fitting for the amplitude of this Gaussian function, $Q_{\rm BF}$ and $U_{\rm BF}$. We also derive from this fit the errors $\sigma(Q_{\rm BF})$ and $\sigma(U_{\rm BF})$.
Polarized intensity, $P$, maps in Fig. 6 are calculated as $\sqrt{Q^{2}+U^{2}}$ for visualization purposes. However, the estimates for $P$ used to compute the polarization efficiency are not directly derived from the $P$ maps; rather, they come from the $Q$ and $U$ BF estimates. We divide the fitted $P$ by the fitted $I$ and compare the result with the model. This approach minimizes the noise bias that affects polarized intensity measurements in regions with low signal-to-noise ratios.
Table 4: Raster scan observation properties and characterization of Pixel 23 from Tau A. The upper part lists the scan parameters, while the lower part reports polarization efficiency ($\eta$) and responsivity ($\mathfrak{R}$) for each channel (CH). [^7] Single Tau A observation conducted on 30 November 2021.
@@
We use the ratio of the measured polarization fraction to the expected value from previous measurements to estimate the polarization efficiency of our detectors. We assumed the expected value would be constant and equal to $7\pm 0.1$ % at TFGI frequencies. We obtained this value from the weighted average of the results for the five frequency bands given in Weiland et al. (2011). We did not propagate this uncertainty to the final map measurements, as the errors in $IQU$ measurements dominate. We find an average polarization efficiency across the channels of $\eta=95\pm 3$% (the uncertainty is calculated as the standard deviation across the channels), with a single-channel efficiency between $\eta\sim$91–98 % (as seen in Table 7).

### v2

@@ Abstract
Context. Polarization measurements in the 30–40 GHz range are essential to improve the understanding of Galactic synchrotron radiation and potential anomalous microwave emissions. Ultimately, this advancement will aid in detecting primordial B modes. To accomplish this, it is crucial to achieve unprecedented performance in instrumentation. Specifically, accurately calibrating instrument responses and polarization angle is an essential step in state-of-the-art setups. The modeling and mitigation of systematic effects are critical in this context.
@@
Results. By creating calibration datasets obtained with the reference diode, we evaluate the data quality and quantify phase switch errors to account for the fine polarization response. We also utilize Tau A and Moon observations to calibrate the system’s response and stability over time. In addition, we calculate the refraction index of the Moon to be $n_{\rm{Moon}}=1.209\pm 0.007\,\text{(stat)}\,\pm 0.009\,\text{(sys)}$ at 31 GHz under smooth-surface assumption.
@@ 1 Introduction
With the completion of these groundbreaking observations, attention has shifted towards investigating the polarization of the CMB (Abazajian et al. 2016; Ade et al. 2019; Hazumi et al. 2019). Despite its signal being orders of magnitude fainter than total intensity, advancements in instrumentation now permit the study of polar signals with unprecedented sensitivity. The polarization pattern of the CMB, typically decomposed into E-modes and B-modes (Seljak & Zaldarriaga 1997; Kamionkowski et al. 1997), carries critical information about the early Universe’s conditions, including the presence of primordial gravitational waves generated during inflation.While satellite-based experiments are required to measure the largest angular scales, ground-based observations have also made significant contributions to the field, especially at small angular scales where satellite telescopes struggle to achieve the necessary angular resolution. The international thrust has deployed in the last decade, among others, the Atacama B-mode Search experiment (ABS; Simon et al. 2016; Kusaka et al. 2018), the Atacama Cosmology Telescope (ACT; Li et al. 2021; Qu et al. 2024), the BICEP/Keck Array program (Hui et al. 2018; BICEP/Keck Collaboration et al. 2022), the Cosmology Large Angular Scale Surveyor (CLASS; Essinger-Hileman et al. 2014; Eimer et al. 2024), the GroundBIRD telescope (Tsujii et al. 2024), the Polarbear experiment (Inoue et al. 2016; Adachi et al. 2022, the predecessor of the Simons Array experiment; Stebor et al. 2016), the South Pole Telescope (SPT; Chown et al. 2018; Sobrin et al. 2022), Spider (Ade et al. 2022; Shaw et al. 2024), and the Q & U Bolometric Interferometer for Cosmology (QUBIC; Mennella et al. 2024).
In this context, the Q-U-I JOint TEnerife (QUIJOTE) CMB experiment aims to achieve two primary scientific objectives: firstly, to measure the tensor-to-scalar ratio (r) if larger than $\sim$0.05 (Rubiño-Martín et al. 2017); and secondly, to characterize polarized foregrounds (Betoule et al. 2009), particularly at low frequencies, shedding light on the synchrotron and anomalous microwave emissions. The current best upper limit on $r$ is $r<0.032$, at a 95% confidence level (Tristram et al. 2022). Current-generation experiments like the Simons Observatory small aperture telescopes (SO-SATs) aim at constraining $r$ with $\sigma(r)=0.003$ or better (Ade et al. 2019), while next-generation telescopes like LiteBIRD will attempt a $r<0.002$, at a 95% confidence level (LiteBIRD Collaboration et al. 2023).
@@
In addition to polarization angle calibration, an equally crucial aspect is the precise calibration of detector responsivity. Any uncertainty in the responsivity directly affects the amplitude of the measured Stokes parameters, thereby introducing a multiplicative bias in the polarization power spectra. For next-generation CMB experiments aiming to detect primordial B-modes at levels of $r<10^{-3}$, the required responsivity calibration must be at the sub-percent level, typically better than 0.5%, to avoid systematic errors that could bias the signal (Abitbol et al. 2021). Responsivity mismatches across detectors can also induce leakage from temperature anisotropies into polarization. Precise calibration between frequency channels is also required to ensure a reliable component separation and then minimize contamination on the measurement of the cosmological signal from foreground residuals (Carralot et al. 2025). Finally, differential responsivity errors distort the reconstructed Stokes parameters by introducing mixing between $I$, $Q$, and $U$, which in turn leads to leakage between $T$, $E,$ and $B$ (O’Dea et al. 2007).
In this work, we describe the calibration of the TFGI system by characterizing its central detector, Pixel 23 at 31 GHz. This detector is the one located in the center of the focal plane, and we use it as a benchmark for the full focal-plane calibration. This choice is driven by the central pixel’s optimal optical performance; however, the results are not assumed to be directly representative of off-axis detectors. A full focal-plane calibration is beyond the scope of this work, which primarily aims to establish a calibration methodology for TFGI. In Sect. 2, we provide an overview of the TFGI instrument. Section 3 details the ground measurement performed with the instrument mounted on the telescope but without sky scanning, including its specific setup and mathematical framework (Sect. 3.1), data structure (Sect. 3.2) for polarization efficiency (Sect. 3.3), responsivity (Sect. 3.4), polarization angle (Sect. 3.5), and phase error (Sect. 3.6) characterization. Section 4 presents the results from an on-sky calibration and response stability using Tau A (Sect. 4.1) and the Moon (Sect. 4.2). Section 5 provides a comparison of the results across different setups. In addition, we calculate the refraction index of the Moon in Appendix A. Finally, the conclusions are presented in Sect. 6.
@@ 2 The QUIJOTE-TFGI instrument
TFGI was first assembled with 29 receivers in 2016 and achieved its technical first light in May 2016 (Pérez-de-Taoro et al. 2016; Rubiño-Martín et al. 2017), but observations were interrupted due to cryostat leak issues. After preliminary commissioning phases carried out between 2018 and 2019, the TFGI was finally installed at the second QUIJOTE telescope (QT2) and underwent its main commissioning phase from November 2021 to October 2022 with only 7 pixels, 4 at 31 GHz and 3 at 41 GHz. Scientific results from this commissioning phase will be presented in a separate paper (Fernández-Torreiro, M. et al. in prep.). The data analyzed in this work correspond exclusively to this latter commissioning period.
@@ 3.1 Setup
The calibration process involves positioning the so-called multi-frequency calibration instrument (MFCI) in front of the detector and acquiring calibration data, as illustrated in Fig. 2. The MFCI consists of two feedhorns and OMTs, similar to the cryogenically cooled feedhorns and OMTs of the TFGI, and is connected back-to-back through two identical -20 dB waveguide couplers. The calibration signal is introduced through one coupler, while the other is terminated with a room-temperature load. The horns are aligned through a concentric stepper motor axial drive, and this setup is designed to be attached to a detector-centered polarization grid known as a fixed plate. Additionally, a tilted aluminum flat mirror points the horn arrangement toward the sky when the cryostat is mounted on the focal plane of the QT2. In this way, the background radiation during a calibration run comes from the sky. The injected signal corresponds to the diode’s equivalent temperature when ON with a 20 dB coupling ratio. A wideband noise diode (Keysight 346CK01111https://www.keysight.com/us/en/product/346CK01/noise-source-1-ghz-50-ghz.html) was selected for its coverage of both FGI and TGI bands.
For installation, the MFCI is preassembled horizontally using a digital inclinometer and mounted onto the telescope focal plane with 0.2 deg mechanical repeatability of the MFCI mounting angle when reinstalling the calibration plate. A dedicated acquisition system controls the stepper motor, the diode, and the phase switches. Polarization angles are referenced with respect to the diode’s E-field and defined as positive for anticlockwise rotation from the sky toward the instrument. The alignment plate is secured over the cryostat focal plane, and the MFCI is attached with spring-loaded clamps, ensuring the mirror maintains its skyward orientation. Electrical connections for the motor, diode, and switches are routed through dedicated cabling.
@@ 3.1.1 Process of polarization characterization and mathematical framework
where V is the output in volts [^2] Not to be mistaken this time with the Stokes $V$ parameter., $g$ is the gain, $\eta$ is the polarization efficiency, $\delta$ is the phase state of each polarization state (0, 90, 180, and 270 deg), and $\epsilon$ is the correction angle that represents the deviation from the nominal phase value assumed constant over the band. The subscript indicates the channel, with a comma distinguishing between TGI and FGI. Also $g$, $\epsilon$, and $\phi_{\rm{c}}$ differ between TGI and FGI. Operationally, we use Eq. (2c) for all channels in the on-sky analysis, since the four-channel equations are mathematically equivalent and differ only by successive 45-degree rotations; adopting a single convention therefore simplifies the on-sky data processing without loss of generality.
@@ 3.2.1 Data assessment with diode reference
Figure 4: The histogram of the slopes that quantify the signal stability in a data acquisition of 1 minute. Each count is defined as a function of the slope obtained from linear fitting. The slope values span a range $\sim$2.5-2.5–0.5 mV/minute for the diode ON case, indicating a lack of significant stability issues within 1 minute of integration.
The data present a linear regression slope of a few hundred mV/minute; $1.6\pm 0.8$ mV/minute in the diode ON case and $2.5\pm 0.3$ mV/minute in the diode OFF case. The uncertainties are calculated as the standard deviation of the counts. This leads to a drift in the signal smaller than 2 mV in a minute. Equivalent to $\sim$50 mK by applying the average on-sky responsivity reported in Sect. 5.2. This instability is likely due to a combination of gain drift and the instability of the optical load, originating from the background and the diode. Although this small drift is detectable, it has a negligible impact on the results discussed here. Firstly, it is comparable to the white-noise level. Secondly, it is subdominant with respect to other systematic effects in the calibration. In particular, differences between the four phase-state configurations, which should nominally produce identical values, dominate the uncertainty budget at the few-percent level, whereas the measured drift over one minute is only at the millipercent level.
@@ 3.4 Responsivity calibration
The responsivity calculated from the diode is primarily affected by the uncertainty in the diode’s emitted power as seen by the detector, and the fact that this measurement does not reflect the actual observational configuration used on-sky – for instance, it does not include coupling through the QT2 mirrors, which alters the beam pattern and effective system gain. While the detector’s intrinsic responsivity should, in principle, be independent of the optical setup, the total optical coupling efficiency and beam gain introduced by the optics do influence the system-level responsivity. Additionally, the relatively high temperature of the diode may induce nonlinearities in the detector response, further complicating the calibration. The calibration diode is mainly used to monitor system responsivity stability rather than to provide an absolute responsivity calibration. However, future measurements will incorporate active thermal control and a lower-output diode to further mitigate possible non-linear effects.
@@ 3.6 Phase errors calibration
The initial fitting step aims to constrain the $\phi_{c}$ parameter for each channel. We fit Eq. (2c) under the assumption of $\epsilon=0$ deg. Consequently, the resulting $\phi_{c}$ values encompass the average $\epsilon$ across different phase angles. Subsequently, we perform the fitting for each dataset, with 19, 15, and 15 iterations, respectively, for the three datasets. We exploit the results reported in Table 1 (last row) as the initial guess of the fitting for $\phi_{c}$. The results are reported in Table 5.
@@
Table 3: Phase error corrections ($\epsilon$). [^6] The results are presented for each channel (CH) and phase state ($\delta$). The values represent averages across the datasets used, with the uncertainties calculated as the standard deviation among the different datasets. The final row reports the mean values for each phase state, with uncertainties obtained by propagating the uncertainties in the mean calculation across channels, based on the expectation that the errors should be consistent across channels by design. We observe that the quoted uncertainties come from only three datasets tied to this specific instrument setup. Although this limits the statistical strength of the estimates, it reflects the available commissioning data and offers a useful indication of parameter stability; future calibration campaigns will aim to improve this aspect.
@@
[…]
@@ 4.1.1 Polarization efficiency with Tau A
For this analysis, we selected the Tau A observation with the best noise properties while being close in time to the Moon observation that will be discussed in Sect. 4.2. This observation consists of a single raster scan, and its properties are summarized in Table 4. From this observation, we produce maps of the three Stokes parameters for each channel independently, which are shown in Fig. 6. The detection of a polarized signal from Tau A is immediately clear. To estimate the source flux density, we use a beam-fitting (BF) photometry that involves fitting the $I$, $Q$, and $U$ maps of each channel to a 2D Gaussian with fixed position (using the nominal coordinates of Tau A) and width fixed to the FWHM of the beam calculated in Fernández-Torreiro, M. et al. (in prep.). The amplitudes of the Gaussians obtained from these fits define $I_{\rm BF}$, $Q_{\rm BF}$ and $U_{\rm BF}$ estimates, with their uncertainties being $\sigma(I_{\rm BF})$, $\sigma(Q_{\rm BF})$ and $\sigma(U_{\rm BF})$. At the TFGI angular resolution, Tau A is only marginally resolved and is therefore well described by a Gaussian profile; any residual source-extension effects are absorbed into the BF uncertainty. As a consistency check, we also performed aperture photometry on the Tau A observations and obtained results consistent with the BF analysis.
Polarized intensity, $P$, maps in Fig. 6 are calculated as $\sqrt{Q^{2}+U^{2}}$ for visualization purposes. However, the estimates for $P$ used to compute the polarization efficiency are not directly derived from the $P$ maps; rather, they come from the $Q$ and $U$ BF estimates, $Q_{\rm BF}$ and $U_{\rm BF}$. We divide this $P$ by the fitted $I_{\rm BG}$ and compare the result with the model. This approach minimizes the noise bias that affects polarized intensity measurements in regions with low signal-to-noise ratios.
Table 4: Raster scan observation properties and characterization of Pixel 23 from Tau A. The upper part lists the scan parameters, while the lower part reports polarization efficiency ($\eta$) and responsivity ($\mathfrak{R}$) for each channel (CH). [^7] Single Tau A observation conducted on 30 November 2021. We chose this scan as it showed one of the lowest noise estimates of our sample, and it was close in time to the Moon observations from Section 4.2. Additional scans are included in the monthly coadded analysis presented in Table 10 and will be further discussed in Fernández-Torreiro, M. et al. in prep.
@@
We use the ratio of the measured polarization fraction to the expected value from previous measurements to estimate the polarization efficiency of our detectors. We assumed the expected value would be constant and equal to $7\pm 0.1$ % at TFGI frequencies. We obtained this value from the weighted average of the results for the five frequency bands given in Weiland et al. (2011). Although the FWHM values in Weiland et al. (2011) bands are larger than the one from TFGI, the signal from Tau A is fully contained in one beam in both cases. Because of that, we do not expect a change in the polarization fraction, as shown by previous works (see e.g. Fig. 39 from Rubiño-Martín et al. 2023). We did not propagate this 0.1% uncertainty to the final map measurements, as the errors in $IQU$ measurements dominate. We find an average polarization efficiency across the channels of $\eta=95\pm 3$% (the uncertainty is calculated as the standard deviation across the channels), with a single-channel efficiency between $\eta\sim$91–98 % (as seen in Table 4).